The sum of three numbers is 74 . The first number is 10 more than the second . The third number is 2 times the second. What are the numbers?


Answer by rfer(12644)

1=x+10=26
2=x=16
3=2x=32
------------
x+10+x+2x=74
4x=74-10
4x=64
x=64/4
x=16